# Constrained instantons in scalar field theories

### v1

@@ 1 Introduction
A classic example is vacuum decay. It occurs typically in theories where the potential has a local minimum, termed the false vacuum. In classical field theory a local minimum is stable, but in quantum field theory the field is allowed to tunnel from the false vacuum to the true vacuum, i.e., the global minimum. When a corresponding instanton solution exists, its action determines the rate of this transition Coleman:1977py .
@@
Despite these successes, the constrained instanton approach has so far been limited to cases in which the constrained instantons are small perturbations of the instanton solutions of the massless theory [^1] The authors are aware of one attempt at a non-perturbative construction of a constrained instanton in $SU(2)$ Yang-Mills-Higgs theory — see refs. Klinkhamer:1993kn ; Klinkhamer:1996ad .. This is manifested in the use of the instanton “size” – a concept restricted to the massless theory – to determine the relevant solution, and in the use of perturbation theory to find the solution. In this paper we present a fully non-perturbative formulation of the idea, which allows us to find the relevant solutions even when they are not well approximated by massless instantons. We also obtain an explicit expression for the vacuum decay rate in terms of these constrained instanton solutions.
We begin section 2 with a review of the physics of vacuum decay in theories with exact instanton solutions. We then show how constrained instantons can be used to compute the decay rate in theories with no instantons, and how the use of Lagrange multipliers helps simplify the calculation. We also derive an analytic expression for the constrained instanton rate. In section 3 we apply our improved approach to the case of a massive scalar field with negative quartic term in the potential and we discuss the allowed forms of the constraint. We also describe a method of counting the number of negative modes in the spectrum of a given solution in the presence of a constraint, which is crucial to determine whether it is in fact a constrained instanton, and therefore whether it contributes to the decay rate. We find the constrained solutions and their action numerically in section 4 for two different types of constraint, allowing us to highlight the similarities and differences due to two reasonable but distinct choices of constraint. Finally, we conclude in section 5.
@@ 2.1 Vacuum decay rate
In order to compute eq. (2), we are interested in stationary points that would give an imaginary contribution. This means that the fluctuation determinant has to be negative, and therefore the fluctuation operator (4) has to have an odd number of negative modes. In practice, the dominant contribution usually comes from instantons — saddle point configurations with a single negative mode which are localised in spacetime. [^2] Topological instantons such as the Yang-Mills instanton are an exception, as they do not have negative modes. Such a configuration breaks translation invariance, which means that it has one zero mode per spacetime dimension. Integration over these zero modes using collective coordinates gives a spacetime volume factor, so by defining $\overline{\text{Det}}$ as the functional determinant with these zero modes removed, we have
$\int{\cal D}\phi e^{-S_{E}[\phi]}\approx Z[\phi^{0}]\left[1+i{\cal V}\left|\frac{{\overline{\text{Det}}}\,M[\phi^{\rm inst}]}{\text{Det}\,M[\phi^{0}]}\right|^{-1/2}e^{-\left(S_{E}[\phi^{\rm inst}]-S_{E}[\phi^{0}]\right)}\right]~,$ (5)
where $\phi^{\rm inst}$ is the instanton configuration. Substituting this into eq. (2) gives
$\Gamma\approx 2\left|\frac{{\overline{\text{Det}}}\,M[\phi^{\rm inst}]}{\text{Det}\,M[\phi^{0}]}\right|^{-1/2}e^{-\left(S_{E}[\phi^{\rm inst}]-S_{E}[\phi^{0}]\right)}~.$ (6)
@@
$\Gamma\approx 2\int d\bar{\xi}\,\left|\frac{{\overline{\text{Det}}}_{\bar{\xi}}M_{\bar{\xi}}[\phi^{\rm inst}_{\bar{\xi}}]}{{\text{Det}}M[\phi^{0}]}\right|^{-1/2}e^{-\left(S_{E}[\phi^{\rm inst}_{\bar{\xi}}]-S_{E}[\phi^{0}]\right)}~.$ (12)
@@ 2.2 Lagrange multiplier
$\int dy\,\tilde{M}_{\kappa}(x,y)\psi(y)=\zeta(x)~.$ (21)
@@
$\Gamma\approx 2\int d\bar{\xi}\,\left|\frac{\nu(\bar{\xi})\,{\overline{\text{Det}}}\tilde{M}_{\kappa}[\phi^{\rm inst}_{\bar{\xi}}]}{{\text{Det}}M[\phi^{0}]}\right|^{-1/2}e^{-\left(S_{E}[\phi^{\rm inst}_{\bar{\xi}}]-S_{E}[\phi^{0}]\right)}~,$ (22)
@@ 4.2 $\phi^{3}$ constraint
To interpret the two branches of solutions, we studied the number of negative modes in the constrained fluctuation spectrum around the solutions for all investigated values of $\kappa$, following the procedure described in section 3.3. Unsurprisingly, we found that there is precisely one negative eigenvalue in the spectrum of the unconstrained operator $\tilde{M}_{\kappa}[\phi]$ (eq. (18)) for all values of $\kappa$. Therefore, depending on the sign of the projection factor $\nu$ (eq. (19)), the constrained spectrum can contain up to one negative eigenvalue.
@@ 5 Discussion and conclusions
Focusing on the constrained instanton solutions (i.e. the upper branch) we can see that the constrained instanton action is everywhere higher than that of the massless instanton, and it increases but remains finite as $\bar{\xi}$ approaches the end of the allowed range, with a different maximum value for the two constraints. It is generally well approximated by the massless instanton action, and the actual solutions shown in figures 2 and 8 also resemble the massless instanton solutions for small $|\kappa|$, which suggests that some kind of perturbative approach should be reasonably accurate. We will study this in more detail in a future work asymptotic-limits .
The vacuum decay rate can be obtained from eq. (22) by integrating over the constraint $\bar{\xi}$, but the numerical calculation is beyond the scope of this paper. As a physical quantity, the rate should be independent of the choice of the constraint function, and therefore comparing the results obtained using the two constraints will provide a good indicator of the accuracy of the approximations involved in our method.
@@ Appendix A Constrained functional determinant
With this choice of coordinates, $M_{\bar{\xi}}$ is then simply the submatrix of $\tilde{M}_{\kappa}$ obtained by deleting the $N$th row and the $N$th column, and its determinant is the corresponding cofactor $\tilde{m}_{NN}$ of $\tilde{M}_{\kappa}$. A standard linear algebra result allows us to express the inverse of $\tilde{M}_{\kappa}$ in terms of its determinant and cofactors $\tilde{m}_{ij}$ as
@@
$\int dy\tilde{M}_{\kappa}(x,y)\psi(y)=\zeta(x)~,$ (76)
@@ Appendix B Constrained instanton solutions — consistency checks
There exist multiple ways of checking whether the field configurations we found are indeed the constrained instantons we were looking for. These checks make use of the properties of Lagrange multipliers described in section 2.2. We performed three separate checks on our solutions.
@@ Appendix C Numerics — consistency checks
We were also able to determine the exponents and prefactors, $a$, $b$, $c$, and $d$. These are shown in tables 1 and 2 . It seems that for both constraints and all values of $\kappa$, the value of $b$ is consistent with $2$, and the value of $d$ is consistent with $4$. The prefactors $a$ and $c$ are all different, which is not surprising — it is expected that the prefactors will be heavily dependent on the corresponding instanton size, and therefore on the value of $\kappa$.

### v2

@@ 1 Introduction
A classic example is vacuum decay. It occurs typically in theories where the potential has a local minimum, termed the false vacuum. In classical field theory a local minimum is stable, but in quantum field theory the field is allowed to tunnel from the false vacuum to the true vacuum, i.e., the global minimum. When a corresponding instanton solution exists, its action determines the rate of this transition Coleman:1977py ; Callan:1977pt .
@@
Despite these successes, the constrained instanton approach has so far been limited to cases in which the constrained instantons are small perturbations of the instanton solutions of the massless theory [^1] The authors are aware of one attempt at a non-perturbative construction of a constrained instanton in $SU(2)$ Yang-Mills-Higgs theory — see refs. Klinkhamer:1993kn ; Klinkhamer:1996ad .. This is manifested in the use of the instanton “size” – a concept restricted to the massless theory – to determine the relevant solution, and in the use of perturbation theory to find the solution. In this paper we present a fully non-perturbative formulation of the idea, which allows us to find the relevant solutions even when they are not well approximated by massless instantons. We also obtain an explicit mathematical expression for the vacuum decay rate in terms of these constrained instanton solutions in a generic scalar field theory. We apply our method to a simple toy model, in which we obtain the constrained instanton solutions.
We begin section 2 with a review of the physics of vacuum decay in theories with exact instanton solutions. We then show how constrained instantons can be used to compute the decay rate in theories with no instantons, and how the use of Lagrange multipliers helps simplify the calculation. We also derive an analytic expression for the constrained instanton rate. In section 3 we apply our improved approach to the case of a massive scalar field with negative quartic term in the potential and we discuss the allowed forms of the constraint. We also describe a method of counting the number of negative modes in the spectrum of a given solution in the presence of a constraint, which is crucial to determine whether it is in fact a constrained instanton, and therefore whether it contributes to the decay rate. We find the constrained solutions and their action numerically in section 4 for two different types of constraint, allowing us to highlight the similarities and differences due to two reasonable but distinct choices of constraint, but we leave the numerical computation of the full decay rate for future work. Finally, we conclude in section 5.
@@ 2.1 Vacuum decay rate
In order to compute eq. (2), we are interested in stationary points that would give an imaginary contribution. This means that the fluctuation determinant has to be negative, and therefore the fluctuation operator (4) has to have an odd number of negative modes. In practice, the dominant contribution usually comes from instantons — saddle point configurations with a single negative mode which are localised in spacetime. [^2] Topological instantons such as the Yang-Mills instanton are an exception, as they do not have negative modes. Such a configuration breaks translation invariance, which means that it has one zero mode per spacetime dimension. Integration over these zero modes using collective coordinates gives a factor proportional to the spacetime volume, so by defining $\overline{\text{Det}}$ as the functional determinant with these zero modes removed, we have Callan:1977pt
$\int{\cal D}\phi e^{-S_{E}[\phi]}\approx Z[\phi^{0}]\left[1+i\frac{{\cal V}}{2}\,\left(\frac{S[\phi^{\rm inst}]}{2\pi}\right)^{2}\left|\frac{{\overline{\text{Det}}}\,M[\phi^{\rm inst}]}{\text{Det}\,M[\phi^{0}]}\right|^{-1/2}e^{-\left(S_{E}[\phi^{\rm inst}]-S_{E}[\phi^{0}]\right)}\right]~,$ (5)
Substituting this into eq. (2) gives
$\Gamma\approx{}\left(\frac{S[\phi^{\rm inst}]}{2\pi}\right)^{2}\left|\frac{{\overline{\text{Det}}}\,M[\phi^{\rm inst}]}{\text{Det}\,M[\phi^{0}]}\right|^{-1/2}e^{-\left(S_{E}[\phi^{\rm inst}]-S_{E}[\phi^{0}]\right)}~.$ (6)
@@
$\Gamma\approx{}\int d\bar{\xi}\,\left(\frac{S[\phi^{\rm inst}_{\bar{\xi}}]}{2\pi}\right)^{2}\left|\frac{{\overline{\text{Det}}}_{\bar{\xi}}M_{\bar{\xi}}[\phi^{\rm inst}_{\bar{\xi}}]}{{\text{Det}}M[\phi^{0}]}\right|^{-1/2}e^{-\left(S_{E}[\phi^{\rm inst}_{\bar{\xi}}]-S_{E}[\phi^{0}]\right)}~.$ (12)
@@ 2.2 Lagrange multiplier
$\int d^{4}y\,\tilde{M}_{\kappa}(x,y)\psi(y)=\zeta(x)~.$ (21)
@@
$\Gamma\approx{}\int d\bar{\xi}\,\left(\frac{S[\phi^{\rm inst}_{\bar{\xi}}]}{2\pi}\right)^{2}\left|\frac{\nu(\bar{\xi})\,{\overline{\text{Det}}}\tilde{M}_{\kappa}[\phi^{\rm inst}_{\bar{\xi}}]}{{\text{Det}}M[\phi^{0}]}\right|^{-1/2}e^{-\left(S_{E}[\phi^{\rm inst}_{\bar{\xi}}]-S_{E}[\phi^{0}]\right)}~,$ (22)
@@ 4.2 $\phi^{3}$ constraint
To interpret the two branches of solutions, we studied the number of negative modes in the constrained fluctuation spectrum around the solutions for all investigated values of $\kappa$, following the procedure described in section $3.$ Unsurprisingly, we found that there is precisely one negative eigenvalue in the spectrum of the unconstrained operator $\tilde{M}_{\kappa}[\phi]$ (eq. (18)) for all values of $\kappa$. Therefore, depending on the sign of the projection factor $\nu$ (eq. (19)), the constrained spectrum can contain up to one negative eigenvalue.
@@ 5 Discussion and conclusions
Focusing on the constrained instanton solutions (i.e. the upper branch) we can see that the constrained instanton action is everywhere higher than that of the massless instanton. It seems to approach the massless instanton action in the limit $\bar{\xi}\to 0$ for the $\phi^{3}$ constraint and $\bar{\xi}\to\infty$ for the $\phi^{6}$ constraint, which both correspond to the $\kappa\to 0$ limit. It then monotonically increases while remaining finite as $\bar{\xi}$ approaches the end of the allowed range, with a different maximum value for the two constraints. It is generally well approximated by the massless instanton action, and the actual solutions shown in figures 2 and 8 also resemble the massless instanton solutions for small $|\kappa|$, which suggests that some kind of perturbative approach should be reasonably accurate. We will study this in more detail in a future work asymptotic-limits .
The vacuum decay rate can be obtained from eq. (22) by integrating over the constraint $\bar{\xi}$.In the limit where the action approaches the massless instanton action, the exponential factor in the integrand is largest. However, in order to compute the rate accurately, one must calculate the functional determinants. This calculation and therefore the integration has to be carried out numerically and is left for future work. As a physical quantity, the rate should be independent of the choice of the constraint function, and therefore comparing the results obtained using the two constraints will provide a good indicator of the accuracy of the approximations involved in our method.
@@ Appendix A Constrained functional determinant
With this choice of coordinates, $M_{\bar{\xi}}$ is then simply the submatrix of $\tilde{M}_{\kappa}$ obtained by deleting the NNth row and the NNth column, and its determinant is the corresponding cofactor $\tilde{m}_{NN}$ of $\tilde{M}_{\kappa}$. A standard linear algebra result allows us to express the inverse of $\tilde{M}_{\kappa}$ in terms of its determinant and cofactors $\tilde{m}_{ij}$ as
@@
$\int d^{4}y\,\tilde{M}_{\kappa}(x,y)\psi(y)=\zeta(x)~,$ (76)
@@ Appendix B Constrained instanton solutions — consistency checks
There exist multiple ways of checking whether the field configurations we found are indeed the constrained instantons we were looking for. These checks make use of the properties of Lagrange multipliers described in section $2.$ We performed three separate checks on our solutions.
@@ Appendix C Numerics — consistency checks
We were also able to determine the exponents and prefactors, $a$, $b$, $c$, and $d$. These are shown in tables 1 and 2 . It seems that for both constraints and all values of $\kappa$, the value of $b$ is consistent with 22, and the value of $d$ is consistent with 44. The prefactors $a$ and $c$ are all different, which is not surprising — it is expected that the prefactors will be heavily dependent on the corresponding instanton size, and therefore on the value of $\kappa$.